OpenStudy (algtrigcalc):

How do you find the center or |z-1|=1? |Z|=|x+yi|=sqrt(x^2+y^2) The answer is suppose to be (1,0)

satellite73 (satellite73):

that says the distance between \(z\) and 1 is 1

satellite73 (satellite73):



satellite73 (satellite73):

all points in the complex plane whose distance from 1 is 1 unit , so a circle 

satellite73 (satellite73):

you need some kind of algebra method?

satellite73 (satellite73):

if \(z=a+bi\) then \(z-1=(a-1)+bi\) and so \[|(a-1)+bi|=\sqrt{(a-1)^2+b^2}=1\]

OpenStudy (algtrigcalc):

Oh I see, thank you so much!

satellite73 (satellite73):

or in your notation \[(x-1)^2+y^2=1\] a circle with center \((1,0)\) and radius 1 yw